You are evaluating a project for The Ultimate recreational tennis racket, guaranteed to correct that wimpy backhand. You estimate the sales price of The Ultimate to be $440 per unit and sales volume to be 1,000 units in year 1; 1,250 units in year 2; and 1,325 units in year 3. The project has a 3-year life. Variable costs amount to $245 per unit and fixed costs are $100,000 per year. The project requires an initial investment of $177,000 in assets, which will be depreciated straight-line to zero over the 3-year project life. The actual market value of these assets at the end of year 3 is expected to be $39,000. NWC requirements at the beginning of each year will be approximately 20 percent of the projected sales during the coming year. The tax rate is 35 percent and the required return on the project is 11 percent. (Use SL depreciation table)


What will the cash flows for this project be? (Negative amounts should be indicated by a minus sign. Do not round intermediate calculations. Round your final answers to 2 decimal places.)
Answer #1

Particulars Year 1 Year 2 Year 3 Year 0 (177,000) (88,000) (22,000) 126,750 20,650 (6,600) 158,438 20,650 Upfront Investment

Investment in NWC will be released at the end of the project.

NPV = Discounted Cash inflows - Discounted Cash outflows

=494658 - 265000 = 229,658

Negative amounts are in closed bracket, you can replace it with minus sign.

Working:

Year 3 1,250 440 Sales in Units (a) Sales price per unit (b) Variable cost per unit (C) Sales (a) * (b) Less: Variable cost (

Saving in tax due to depreciation Cost of Asset 177,000 Useful life Depreciation Per year (x) 59,000 Tax Rate (y) 35% Tax Sav

Discount Factor Formula:

Discount Factor = 7 (1+i)n

Where, i = discount rate

n = number of periods